OpenStudy (lindsey11x16):

A 4.98 g sample of aniline (C6H5NH2, molar mass = 93.13 g/mol) was combusted in a bomb calorimeter with a heat capacity of 4.25 kJ /°C. If the temperature rose from 29.5 ° C to 69.8 ° C, determine the value of ΔH° comb for aniline in kJ/mol. I don't know what °Hcomb even means...my textbook doesn't help either. :( i don't know what formula to use or where to even begin! Can anyone help? this problem is exhausting..

OpenStudy (matt101):

For any reaction you have, there is an associated change in enthalpy (represented by ΔH). You can think of enthalpy as the change in "heat" due to the reaction. So reactions with a positive ΔH gain heat (are endothermic), while those with a negative ΔH lose heat (are exothermic). Sometimes, to be more precise, we name the ΔH to indicate what reaction we're talking about. ΔH comb refers to the heat of combustion, the enthalpy change coming from the combustion reaction that occurs when you burn aniline. ΔH(f) would be the heat of formation of aniline, or the change in enthalpy arising from the reaction that produces aniline from it's constituent elements. There is ΔH(vap) (heat of vaporization), ΔH(fus) (heat of fusion or melting), and more. It's all the same thing, the extra bit just specifies the reaction we're talking about. Now back to your question...let's not get exhausted! Remember the equation ΔHn=mcΔT? What would be your first step?

OpenStudy (thadyoung):

\(\Delta H_{\sf comb}\) means the change in enthalpy of combustion. That means you have some organic molecule in air and some source of fire, forming carbon dioxide and water

OpenStudy (lindsey11x16):

This is what I know what I have so far for that formula you provided. m = 4.98 g ΔT = 69.8-29.5 = 40.3 C = 4.25 kJ/C (i'm guessing i use the specific heat of aniline) n = do I use the molar mass ? 93.13 g/mol now i'm stuck...again :(

OpenStudy (matt101):

n is referring to the moles of aniline. How do you find moles if you're given mass and molar mass? PS. everything else looks good!

OpenStudy (lindsey11x16):

I found its moles to be 0.05347 mol. so when i plug that in, am I trying to get ΔH by itself?

OpenStudy (matt101):

Yup! I just need to update one thing I said earlier now that I've reread the question. The equation we should be using here is ΔHn=CΔT. Whenever you're using these equations, we're always talking about transfers of heat (each side is equal to q, which is heat energy). Since it's a transfer of heat, one side is where the heat was transferred TO, the other is where the heat was transferred FROM. In this case we want to find ΔH comb of aniline. Since aniline is being burned, that's where the heat is coming from, which means the left side of the equation (ΔHn) is all talking about aniline. That means the right side of the equation is where the heat was transferred to, which in this case is the bomb calorimeter. Notice the units of the heat capacity given: kJ/°C. That means all we need to do is multiply by temperature (°C) to get heat energy (kJ) on it's own. There's no room for mass on this side of the equation. Does that make sense? Looks like you're on the right track now!

OpenStudy (lindsey11x16):

\[ΔH (0.05347)=(4.25)(40.3)\] \[ΔH(0.05347)=171.275\] \[ΔH=171.275/0.05347\] \[ΔH=3203.2 \] so when i put that answer in scientific notation it gives me 3.20 x 10^3. This is close to one of the multiple choice answers given but their answer is Negative while i got positive. Does this have something to do with losing heat and gaining heat part?

OpenStudy (matt101):

Yes! One more thing I should add (sorry it's late over here) is that the equation should be ΔHn=-CΔT. The negative because what is lost by one side is gained by the other (e.g. -3200 kJ on the left means that was lost, but the right side gained 3200 kJ so we need to multiply by -1 to make it positive)! Is that clear now?

OpenStudy (matt101):

And you know ΔH comb is negative from the start because combustion reaction are always going to release heat (be exothermic).

OpenStudy (lindsey11x16):

Thank you for the very detailed explanation. i really needed it to be broken down like this. goodnight (it's late over here too! Lol)

OpenStudy (matt101):

Any time!